Sally travels by car from one city to another. She drives for 29.0 min at 63.0 km/h, 37.0 min at 40.0 km/h, and 45.0 min at 74.0 km/h, and she spends 14.0 min eating lunch and buying gas.
(a) Determine the average speed for the trip.
km/h
(b) Determine the total distance traveled.
km
1
Expert's answer
2013-01-30T11:05:04-0500
V - an average speed
S - a total distance
t - a total time

V = S/t
S = V1*t1 + V2*t2 + V3*t3
t = t1 + t2 + t3 + t4

b)
S = (29/60)h*63km/h + (37/60)h*40km/h +(45/60)h*74km/h =& 110.6 km
a)
V = S/t = 110.6 /( (29 + 37 + 45 + 14)/60 h) = 53.1 km/h

Answer: a)& 53.1 km/h b) 110.6 km